Mr. Blue is a 45 year old who came into the ED for shortness of breath. He has a medical history of asthma. Upon assessment, you find cyanosis of the lips, a productive cough, temp is 102.1, BP 135/89, HR 110, RR 34, and has rapid and shallow breathing. Breath sounds are diminished bilaterally and you hear coarse rhonchi in the upper lobes. X-ray indicates bilateral pneumonia.

ABG is drawn:

pH 7.42

PCO2 24

HCO3 24

1. What does this ABG indicate? Compensated or Uncompensated? Is this respiratory of metabolic? Is this alkalotic or acidotic?


2. What are 3 nursing interventions you can provide to help Mr. Blue clear his airways?


3. What are 2 medications you expect the physician to order for Mr. Blue?
Answer #1

Normally PH should be 7.35-7.45. PaO2 = 75-100 mm Hg Pacoa : 35-45 mm Hg & HCO3 = 22-26 mEq/L. | ABG = Arterial blood gas In